OpenStudy (bhumi273):

write a linear equation in 2 variable to represent the statement, "4 times the cost of table is equal to 7 times the cost of chair" . from the equation, find the cost of 1 chair if the cost of 1 table is Rs.1050.

OpenStudy (michele_laino):

Please note that if I call with "T" the cost of a table, and with "C" the cost of a chair, what is the requested equation?

OpenStudy (jhannybean):

How much is Rs 1050 in dollars btw?

OpenStudy (bhumi273):

idk Jhannybean

OpenStudy (michele_laino):

this equality: 4T= 7C, can be your equation? @bhumi273

OpenStudy (bhumi273):

ok...

OpenStudy (michele_laino):

why?

OpenStudy (michele_laino):

ok! please substitutte your numerical data into the above equation, what do you get?

OpenStudy (michele_laino):

oops...please substitute...

OpenStudy (michele_laino):

please keep in mind that T = 1050

OpenStudy (bhumi273):

ya thts ok.... now d nxt sum....

OpenStudy (michele_laino):

you should get this: 7 C = 4200

OpenStudy (anonymous):

1$ is about 60 Rs

OpenStudy (michele_laino):

please divide both sides of the above equation by 7 @bhumi273

OpenStudy (michele_laino):

you should get this: C = 4200 /7 =...? please continue

OpenStudy (michele_laino):

what is 4200/7=... please?

OpenStudy (michele_laino):

Is your answer C= 600?

OpenStudy (michele_laino):

Please note that I think that your answer is C = Rs. 600